# 3D Unsupervised Learning by Distilling 2D Open-Vocabulary Segmentation Models for Autonomous Driving

### v1

@@ 3.3.1 Superpixel-Superpoint Generation
Given the point cloud $\mathcal{P}$ and images $\mathcal{I}$ defined in Sec. 3.2, we have generated masks $M_{\mathcal{I}}$ and use $\Gamma_{\text{camera}\leftarrow\text{LiDAR}}$ to map the labels $L_{M}$ of masks to $\mathcal{P}$. We regard the set of pixels with corresponding points in the same mask $m_{r}\in M_{\mathcal{I}}$ as a superpixel $S_{r}^{\text{pixel}}\in\mathcal{S}^{\text{pixel}}$, while the corresponding region of the point cloud as a superpoint $S_{r}^{\text{point}}\in\mathcal{S}^{\text{point}}$, $\mathcal{S}^{\text{pixel}}$ and $\mathcal{S}^{\text{point}}$ establish a bijection and ensuring $|\mathcal{S}^{\text{pixel}}|=|\mathcal{S}^{\text{point}}|=R^{\mathcal{S}}\leq|M_{\mathcal{I}}|=R$. Assuming the point cloud backbone $\mathcal{F}_{\theta_{p}}$ comes with an output head $H_{p}$, we replace $H_{p}$ with a trainable projection head $H_{p}^{*}$, projecting the point cloud feature $f_{p}$ of $p\in\mathcal{P}$ into a $D^{*}$-dimensional space such that $Dim_{f_{M}}=Dim_{f_{T}}=D^{*}$, here $f_{M}\in F_{M}$ and $f_{T}\in F_{T}$ refer to mask features and text features provided by the 2D open-vocabulary semantic segmentation models mentioned in Sec. 3.1. Firstly, we apply average pooling and normalization to each group of pixel features ${F}_{r0}=\{f_{p}|p\in S_{r0}^{\text{point}}\}$ guided by superpoints to extract the superpoint embeddings $f_{r0}^{\text{p}}\in F^{\text{superpoint}}$. Then, we normalize $F_{M}$ as the superpixel embeddings $F^{\text{superpixel}}$ and consider the masks-corresponding text features as $F^{text}$. Finally, we employ a tri-modal contrastive loss to align $F^{\text{superpixel}}-F^{\text{superpoint}}$, $F^{\text{text}}-F^{\text{superpoint}}$.
3.3.2 Tri-Modal Contrastive Loss
@@
$\mathcal{L}_{I-P}=\mathcal{L}_{(F_{\text{superpixel}},F_{\text{superpoint }})}=-\frac{1}{R^{\prime}}\sum_{i=0}^{R^{\prime}}\log\left[\frac{e^{(\langle f_{i}^{I},f_{i}^{p}\rangle/\tau)}}{\sum_{j\neq i}e^{(\langle f_{i}^{I},f_{j}^{p}\rangle/\tau)}+e^{(\langle f_{i}^{I},f_{i}^{p}\rangle/\tau)}}\right],$ (1)
@@
Unlike the superpixel-superpoint contrastive loss, the text-superpoint contrastive loss does not exhibit "self-conflict" on classes of a dataset. However, in TMP, to ensure the uniformity of knowledge, we retained the class dictionary $\mathcal{C}$ as discussed in Sec. 3.2. In UOV-baseline or other downstream tasks, texts of the same class $t_{j_{1}},t_{j_{2}}\in c_{i}$ should be considered as positive samples for each other, so treating the point cloud regions corresponding to $t_{j_{1}},t_{j_{2}}$ as negative examples will inadvertently cause "self-conflict". Therefore, for text $t_{j_{0}}\in c_{i}$, we utilize the text feature’s cosine similarity $\langle f_{t_{j_{0}}},f_{t_{j_{s}}}\rangle$ weighted for other texts in the same class $\{t_{j_{s}}\in c_{i},j_{s}=\{1,\ldots,n_{c_{i}}\}\neq j_{0}\}$ as "semi-positive" samples to compute $\mathcal{L}_{T-P}$:
$\mathcal{L}_{T-P}=\mathcal{L}_{(F_{\text{text }},F_{\text{superpoint}})}=-\frac{1}{R^{\prime}}\sum_{i=0}^{R^{\prime}}\log\left[\frac{e^{(\langle f_{i}^{t},f_{i}^{p}\rangle/\tau)}}{\sum_{t_{j}\neq t_{i}}e^{\left((1-\alpha_{ij})\langle f_{i}^{t},f_{j}^{p}\rangle/\tau\right)}+e^{(\langle f_{i}^{t},f_{i}^{p}\rangle/\tau)}}\right],$ (2)
@@
3.4 Approximate Flat Interaction
Through Sec. 3.2,3.3, we noticed three technical difficulties for annotation-free semantic segmentation that need to be solved: 1) Unprojected point cloud regions caused by differing or occluded fields of view (FoV) among devices. This directly results in the region of the point cloud outside the image FoV remains untrained for long periods, thus the untrained area suffers from serious label confusion. 2) Label noise. This arises from matching errors between cameras and LiDAR, as well as noise inherent in 2D open-vocabulary segmentation models. 3) Category misperception by single-modal model. Point cloud information struggles to distinguish certain categories across different conditions, whereas information-rich images can address this issue perfectly.
Inspired by Point-NN [47], we propose a non-parametric network for Approximate Flat Interaction (AFI). AFI essentially expects points to interact only among points lying on approximate planes, thereby preserving labels of relatively small objects, e.g., pedestrians, and vehicles, within a broad perceptual domain. The process of AFI is formulated as:
@@ 3.4 Approximate Flat Interaction
$L_{\mathcal{P}}^{predict}$ represents the predictions in Sec. 3.2, and $\gamma$ indicates the minimum similarity between the directions when their directional features interact. $L_{\mathcal{P}}^{AFI}$, on the other hand, denotes the point cloud labels predicted by the function $AFI(\cdot)$. Meanwhile, we can choose to assist optimization through the pseudo-labels $L_{\mathcal{P}}^{pseudo}$ generated by 2D open-vocabulary segmentation models. A more detailed description of $AFI(\cdot)$ is stated in Appendix A.
@@
During downsampling, AFI passes the directional features of the sampled center point through layer-wise interactions with neighboring points, and binds the correlation between two points based on 1) whether the two points are relevant in the same direction and 2) the tightness of the relevance between correlated directions. Through four rounds of downsampling, point-to-point interactions construct a network that, apart from points at the junctions, AFI ensures the surfaces formed by points on the same network approximate planes, thus tightly controlling feature interactions among points.
@@
The advantages of AFI are evident. 1) Wide-sensing domain: The perception domain for the point clouds with AFI is wide and possesses strong spatial perception capabilities. 2) Detachability: Not only the entire AFI is detachable, but also the auxiliary module for 2D images within AFI is detachable.The effectiveness of AFI can be referenced in Fig. 5
@@
Table 1: Comparisons of 3D annotation-free semantic segmentation results (mIoU) on nuScenes [8] val set.
@@
CLIP2Scene [9][CVPR’23] 0% SPVCNN [48] 20.80
Chen et al. [11][NeurIPS’23] 0% MinkowskiNet 26.80
OpenScene3D [12][CVPR’23] 0% MinkowskiNet 37.03
UOV(ours)+Mask-clip [15] 0% MinkowskiNet 30.35
UOV(ours)+FC-CLIP [16] 0% MinkowskiNet 43.28
UOV(ours)+CAT-Seg [18] 0% MinkowskiNet 42.83
UOV(ours)+SAN [17] 0% MinkowskiNet 47.73
OpenScene [12][CVPR’23] 0% MinkowskiNet 38.37
UOV(ours)+FC-CLIP [16] 0% MinkowskiNet 43.64
UOV(ours)+CAT-Seg [18] 0% MinkowskiNet 43.64
UOV(ours)+SAN [17] 0% MinkowskiNet 47.89
- 100% MinkowskiNet [7] 74.66
4 Experiments
4.0.1 Datasets
To validate the performance of our model, multiple experiments on two large-scale autonomous driving datasets, nuScenes [8] and SemanticKITTI [49, 50] were conducted, as detailed in Sec. 4.1,4.2. In nuScenes, there are 700 scenes for training, while the validation and test set each consists of 150 scenes, comprising a total of 16 semantic segmentation classes. During pre-training, only the train set was utilized, while we validated using specific scenes separated from the train set. SemanticKITTI has 19 classes, with its 22 sequences partitioned into specific train, validation, and test sets.
4.0.2 Implementation Details
We followed the training paradigm of SLidR [39], employed MinkowskiNet18 [7] as the 3D backbone, and used a linear combination of the cross-entropy and the Lovász loss [51] as training objective in annotation-free and downstream tasks. For 2D open-vocabulary segmentation models, we employed FC-CLIP [16], SAN [17], CAT-Seg [18] for both TMP and annotation-free training, while using MaskCLIP [15] as a control group. The generation of mask features and text features were synchronized with the masks and mask labels. FC-CLIP [16] employed panoptic segmentation, distinguishing different instances on thing-classes. MaskCLIP [15], SAN [17], and CAT-Seg [18] utilized semantic segmentation, not distinguishing instances with the same semantics in both TMP and annotation-free training. Their mask features are selected as the average pool of pixel features in semantically identical regions. In Tri-Modal contrastive Pre-training (TMP) of Sec. 3.3, our network was pre-trained for 40 epochs on 4 V100 GPUs with a batch size of 4, which takes about 80 hours. For annotation-free training in Sec. 3.2 and other downstream tasks, the network was trained for 30 epochs on a single V100 GPU with a batch size of 16, which takes about 18 hours. The temperature coefficient $\tau$ in Eq. 1,2 was set to 0.07, and the optimal results achieved for Eq. 4 when $\alpha_{\text{image}}=\alpha_{\text{text}}=0.5$. In Eq. 5 from Sec. 3.4, the minimum similarity $\gamma$ between the directions when their directional features interact, was set to 0.995, implying that the maximum angular disparity of two interact point features is about 5.7°. In the network structure of AFI, downsampling was performed four times, with the downsampling rate being $1/3$ for the last three times. Additional details about AFI are provided in Appendix A.
Table 2: Comparisons (mIoU) of different pre-training methods pre-trained on nuScenes [8] and fine-tuned on nuScenes and SemanticKITTI [49]. LP denotes linear probing with frozen backbones.
@@ 4.0.2 Implementation Details
PointConstrast(ECCV’20) [36] 21.90 32.50 41.10
DepthConstrast(ICCV’21) [37] 22.10 31.70 41.50
PPKT(arXiv’21) [38] 35.90 37.80 44.00
SLidR(CVPR’22) [39] 38.80 38.30 44.60
ST-SLidR(CVPR’23) [40] 40.48 40.75 44.72
Seal(NeurIPS’23) [41] 44.95 45.84 46.63
UOV-TMP(ours)+FC-CLIP 44.24 45.73 47.02
UOV-TMP(ours)+CAT-Seg 43.95 46.61 48.14
UOV-TMP(ours)+SAN 46.29 47.60 47.72
UOV(ours)+FC-CLIP 52.92 50.58 45.86
UOV(ours)+CAT-Seg 51.02 49.14 47.59
UOV(ours)+SAN 56.35 51.75 46.60
Figure 3: Illustration of image segmentation results of various 2D open-vocabulary segmentation models. We observe that MaskCLIP (pixel-level CLIP) exhibits label confusion and high error rates in semantic segmentation. The output of SEEM not only suffers from missing masks but also contains incorrect mask annotations. The 2D open-vocabulary segmentation results of these two are inferior to those of FC-CLIP, CAT-Seg, and SAN. More 2D open-vocabulary segmentation results are provided in Fig. 6
4.1 Comparison Results
4.1.1 Annotation-free Semantic Segmentation
In Tab. 1, we compare UOV with the most closely related works on 3D semantic segmentation using the unannotated data of nuScenes: CLIP2Scene [9] designs a semantic-driven cross-modal contrastive learning framework. Chen et al. [11] utilizes CLIP and SAM for denoising. OpenScene [12] extracts 3D dense features from an open-vocabulary embedding space using multi-view fusion and 3D convolution, we re-implemented its code using MinkowskiNet18. The optimal result of UOV’s single-modal annotation-free segmentation reaches 47.73% mIoU, surpassing the previous best method by 10.70% mIoU, and the gap between the fully supervised same backbone is only -26.93% mIoU. Under image assistance, it achieves 47.89% mIoU.
4.1.2 Comparisons among 3D Pre-training Methods
@@
We compared the performance of UOV-TMP (only employing TMP) and UOV (employing both steps) against other state-of-the-art methods on multiple downstream tasks in nuScenes and SemanticKITTI, as shown in Tab. 2. All methods utilized MinkowskiNet [7] as the 3D backbone. Most of the compared state-of-the-art methods utilize point cloud-image contrastive learning. SLidR [39] and ST-SLidR [40] employ superpoint-superpixel correspondence granularity, while SEAL [41] employs VLMs to assist in generating superpixels. Our approach achieved optimal results with 1% data fine-tuning on nuScenes and SemanticKITTI, reaching 51.75% mIoU and 48.14% mIoU, respectively, demonstrating a respective improvement of +21.45% mIoU and +8.64% mIoU compared to random initialization. Compared to the previously best results from SEAL, UOV exhibited enhancements of +5.91% mIoU and +1.51% mIoU, respectively. Furthermore, remarkably, the results of the fully supervised linear probing on nuScenes reached 56.35% mIoU, displaying an improvement of +11.40% mIoU compared to SEAL.
4.2 Ablation Study
@@
We conducted a series of ablation experiments on the NuScenes dataset. The ablation targets included different teacher models, TMP, AFI, etc. The results obtained validate the effectiveness of our designs, especially TMP and AFI. Please refer to Appendix B for details of the ablation experiments.
5 Conclusion

### v3

@@ Introduction
Given the point cloud $\mathcal{P}$ and images $\mathcal{I}$, we have generated masks $M_{\mathcal{I}}$ and use $\Gamma_{\text{camera}\leftarrow\text{LiDAR}}$ to map the labels $L_{M}$ of masks to $\mathcal{P}$. We regard the set of pixels with corresponding points in the same mask $m_{r}\in M_{\mathcal{I}}$ as a superpixel $S_{r}^{\text{pixel}}\in\mathcal{S}^{\text{pixel}}$, while the corresponding region of the point cloud as a superpoint $S_{r}^{\text{point}}\in\mathcal{S}^{\text{point}}$, $\mathcal{S}^{\text{pixel}}$ and $\mathcal{S}^{\text{point}}$ establish a bijection and ensuring $|\mathcal{S}^{\text{pixel}}|=|\mathcal{S}^{\text{point}}|=R^{\mathcal{S}}\leq|M_{\mathcal{I}}|=R$. Assuming the point cloud backbone $\mathcal{F}_{\theta_{p}}$ comes with an output head $H_{p}$, we replace $H_{p}$ with a trainable projection head $H_{p}^{*}$, projecting the point cloud feature $f_{p}$ of $p\in\mathcal{P}$ into a $D^{*}$-dimensional space such that $Dim_{f_{M}}=Dim_{f_{T}}=D^{*}$, here $f_{M}\in F_{M}$ and $f_{T}\in F_{T}$ refer to mask features and text features provided by the 2D open-vocabulary semantic segmentation models. Firstly, we apply average pooling and normalization to each group of pixel features ${F}_{r0}=\{f_{p}|p\in S_{r0}^{\text{point}}\}$ guided by superpoints to extract the superpoint embeddings $f_{r0}^{\text{p}}\in F^{\text{superpoint}}$. Then, we normalize $F_{M}$ as the superpixel embeddings $F^{\text{superpixel}}$ and consider the masks-corresponding text features as $F^{text}$. Finally, we employ a tri-modal contrastive loss to align $F^{\text{superpixel}}-F^{\text{superpoint}}$, $F^{\text{text}}-F^{\text{superpoint}}$.
Tri-Modal Contrastive Loss
@@
$\begin{split}&\mathcal{L}_{I-P}=\mathcal{L}{(F_{\text{superpixel}},F_{\text{superpoint }})}\\ &=-\frac{1}{R^{\prime}}\sum_{i=0}^{R^{\prime}}\log\left[\frac{e^{(\langle f_{i}^{I},f_{i}^{p}\rangle/\tau)}}{\sum_{j\neq i}e^{(\langle f_{i}^{I},f_{j}^{p}\rangle/\tau)}+e^{(\langle f_{i}^{I},f_{i}^{p}\rangle/\tau)}}\right],\end{split}$ (1)
@@
Unlike the superpixel-superpoint contrastive loss, the text-superpoint contrastive loss does not exhibit ”self-conflict” on classes of a dataset. However, to ensure the uniformity of knowledge in TMP, we retained the class dictionary $\mathcal{C}$ as discussed in Baseline of AFOV. In downstream tasks, texts of the same class $t_{j_{1}},t_{j_{2}}\in c_{i}$ should be considered as positive samples for each other, so treating the point cloud regions corresponding to $t_{j_{1}},t_{j_{2}}$ as ”negative samples” will inadvertently cause ”self-conflict”. Therefore, for text $t_{j_{0}}\in c_{i}$, we utilize the text feature’s cosine similarity $\langle f_{t_{j_{0}}},f_{t_{j_{s}}}\rangle$ weighted for other texts in the same class $\{t_{j_{s}}\in c_{i},j_{s}=\{1,\ldots,n_{c_{i}}\}\neq j_{0}\}$ as ”semi-positive” samples to compute $\mathcal{L}_{T-P}$:
$\begin{split}&\mathcal{L}_{T-P}=\mathcal{L}{(F_{\text{text }},F_{\text{superpoint}})}\\ &=-\frac{1}{R^{\prime}}\sum_{i=0}^{R^{\prime}}\log\left[\frac{e^{(\langle f_{i}^{t},f_{i}^{p}\rangle/\tau)}}{\sum_{t_{j}\neq t_{i}}e^{\left((1-\alpha_{ij})\langle f_{i}^{t},f_{j}^{p}\rangle/\tau\right)}+e^{(\langle f_{i}^{t},f_{i}^{p}\rangle/\tau)}}\right],\end{split}$ (2)
@@
Approximate Flat Interaction (AFI)
Through the previous two sections of AFOV, we noticed three technical difficulties for annotation-free semantic segmentation that need to be solved: 1) Unprojected point cloud regions caused by differing or occluded fields of view (FoV) among devices. This directly results in the region of the point cloud outside the image FoV remains untrained for long periods, thus the untrained area suffers from serious label confusion. 2) Label noise in 3D. This arises from matching errors between cameras and LiDAR, as well as noise inherent in 2D open-vocabulary segmentation models. Therefore, we need a robust error correction mechanism for AFOV.
Inspired by Point-NN (Zhang et al. 2023), we propose a non-parametric network for Approximate Flat Interaction (AFI). AFI essentially expects points to interact only among points lying on approximate planes, thereby preserving labels of relatively small objects, e.g., pedestrians, and vehicles, within a broad perceptual domain. The process of AFI is formulated as:
@@
$L_{\mathcal{P}}^{predict}$ represents the predictions in Baseline of AFOV, and $\gamma$ indicates the minimum similarity between the directions when their directional features interact. $L_{\mathcal{P}}^{AFI}$, on the other hand, denotes the point cloud labels predicted by the function $AFI(\cdot)$. Meanwhile, we can choose to assist optimization through the pseudo-labels $L_{\mathcal{P}}^{pseudo}$ generated by 2D open-vocabulary segmentation models. A more detailed description of $AFI(\cdot)$ is stated in Appendix A.
@@
During downsampling, AFI passes the directional features of the sampled center point through layer-wise interactions with neighboring points, and binds the correlation between two points based on 1) whether the two points are relevant in the same direction and 2) the tightness of the relevance between correlated directions. Through four rounds of downsampling, point-to-point interactions construct a network that, apart from points at the junctions, AFI ensures the surfaces formed by points on the same network approximate planes, thus tightly controlling interactions among points.
@@
AFI is a robust error correction mechanism for AFOV. The advantages of AFI are evident. 1) Wide-sensing domain: The perception domain for the point clouds with AFI is wide and possesses strong spatial perception capabilities. 2) Detachability: The entire AFI is detachable, and the auxiliary module for 2D images within the AFI is detachable. The effectiveness of AFI can be referenced in Appendix B: Ablation Study.
Experiments
Experiments Setup
Datasets
To validate the performance of our model, multiple experiments on two large-scale autonomous driving datasets, nuScenes (Caesar et al. 2020) and SemanticKITTI (Behley et al. 2019; Geiger, Lenz, and Urtasun 2012) were conducted, as detailed in Comparison Results and Ablation Study. In nuScenes, there are 700 scenes for training, while the validation and test set each consist of 150 scenes, comprising a total of 16 semantic segmentation classes. During pre-training, only the train set was utilized, while we validated using specific scenes separated from the train set. SemanticKITTI has 19 classes, with its 22 sequences partitioned into specific train, validation, and test sets.
Implementation Details
We followed the training paradigm of SLidR (Sautier et al. 2022), employed MinkowskiNet18 (Choy, Gwak, and Savarese 2019) as the 3D backbone, and used a linear combination of the cross-entropy and the Lovász loss (Berman, Triki, and Blaschko 2018) as training objective in annotation-free and downstream tasks. For 2D open-vocabulary segmentation models, we employed FC-CLIP (Yu et al. 2023), SAN (Xu et al. 2023), CAT-Seg (Cho et al. 2023) for both TMP and annotation-free training, while using MaskCLIP (Zhou, Loy, and Dai 2022) as a control group. The generation of mask features and text features were synchronized with the masks and mask labels. FC-CLIP (Yu et al. 2023) employed panoptic segmentation, distinguishing different instances on thing-classes. MaskCLIP (Zhou, Loy, and Dai 2022), SAN (Xu et al. 2023), and CAT-Seg (Cho et al. 2023) utilized semantic segmentation, not distinguishing instances with the same semantics in both TMP and annotation-free training. Their mask features are selected as the average pool of pixel features in semantically identical regions. In Tri-Modal contrastive Pre-training (TMP), our network was pre-trained for 40 epochs on 4 V100 GPUs with a batch size of 4, which takes about 80 hours. For annotation-free training in Baseline of AFOV and other downstream tasks, the network was trained for 5 epochs and 30 epochs on a single V100 GPU, each task taking approximately 3 hours and a batch size of 16. On a 4090 GPU, annotation-free training for 5 epochs only took one hour. The temperature coefficient $\tau$ in Eq. 1,2 was set to 0.07, and the optimal results achieved for Eq. 4 when $\alpha_{\text{image}}=\alpha_{\text{text}}=0.5$. In Eq. 5, the minimum similarity $\gamma$ between the directions when their directional features interact, was set to 0.995, implying that the maximum angular disparity of two interact point features is about 5.7°. In the network structure of AFI, downsampling was performed four times, with the downsampling rate being $1/3$ for the last three times. Additional details about AFI are provided in Appendix A.
@@
Table 1: Comparisons of 3D annotation-free semantic segmentation results (mIoU) on nuScenes (Caesar et al. 2020) val set.
@@
CLIP2Scene [CVPR’23] 0% X SPVCNN 20.80
Chen et al. [NeurIPS’23] 0% X MinkowskiNet 26.80
OpenScene [CVPR’23] 0% X MinkowskiNet 41.30
OpenScene-LSeg 0% X MinkowskiNet 35.50
OV3D [CVPR’24] 0% X MinkowskiNet 44.60
AFOV(ours)+SAN 0% X MinkowskiNet 47.73
OpenScene [CVPR’23] 0% MinkowskiNet 36.30
OpenScene-LSeg 0% MinkowskiNet 42.10
AFOV(ours)+SAN 0% MinkowskiNet 47.89
- 100% MinkowskiNet 74.66
Figure 5: Illustration of image segmentation results of various 2D open-vocabulary segmentation models. We observe that MaskCLIP (pixel-level CLIP) exhibits label confusion and high error rates in semantic segmentation. The output of SEEM not only suffers from missing masks but also contains incorrect mask annotations. More results are provided in Appendix D.
Comparison Results
Annotation-free Semantic Segmentation
In Tab. 1, we compare AFOV with the most closely related works on 3D semantic segmentation using the unannotated data of nuScenes: CLIP2Scene (Chen et al. 2023b) designs a semantic-driven cross-modal contrastive learning framework; Chen et al. (Chen et al. 2023a) utilizes CLIP and SAM for denoising; OpenScene (Peng et al. 2023) extracts 3D dense features from an open-vocabulary embedding space using multi-view fusion and 3D convolution; OV3D (Jiang, Shi, and Schiele 2024) seamlessly aligning 3D point features with entity text features. The optimal result of AFOV’s single-modal annotation-free segmentation reaches 47.73% mIoU, surpassing the previous best method by 3.13% mIoU. Under image assistance, it achieves 47.89% mIoU. The gap between AFOV and the fully supervised same backbone is only -26.93% mIoU.
Table 2: Comparisons (mIoU) of different pre-training methods pre-trained on nuScenes (Caesar et al. 2020) and fine-tuned on nuScenes and SemanticKITTI (Behley et al. 2019). LP denotes linear probing with frozen backbones.
@@
PointConstrast [ECCV’20] 21.90 32.50 41.10
DepthConstrast [ICCV’21] 22.10 31.70 41.50
PPKT [arXiv’21] 35.90 37.80 44.00
SLidR [CVPR’22] 38.80 38.30 44.60
ST-SLidR [CVPR’23] 40.48 40.75 44.72
Seal [NeurIPS’23] 44.95 45.84 46.63
VLM2Scene [AAAI’24] 51.54 47.59 47.37
ScaLR [CVPR’24] 42.40 40.50 -
AFOV-TMP(ours)+FC-CLIP 44.24 45.73 47.02
AFOV-TMP(ours)+CAT-Seg 43.95 46.61 48.14
AFOV-TMP(ours)+SAN 46.29 47.60 47.72
AFOV(ours)+FC-CLIP 52.92 50.58 45.86
AFOV(ours)+CAT-Seg 51.02 49.14 47.59
AFOV(ours)+SAN 56.35 51.75 46.60
Comparisons among 3D Pre-training Methods
@@
We compared the performance of AFOV-TMP (only employing TMP) and AFOV (employing both steps) against other state-of-the-art methods on multiple downstream tasks in nuScenes and SemanticKITTI (all pre-trained on nuScenes), as shown in Tab. 2. All methods utilized MinkowskiNet as the 3D backbone. Most of the compared state-of-the-art methods utilize point cloud-image contrastive learning. SLidR (Sautier et al. 2022) and ST-SLidR (Mahmoud et al. 2023) employ superpoint-superpixel correspondence granularity; ScaLR (Puy et al. 2024) scales the 2D and 3D backbones and pretraining on diverse datasets; while SEAL (Liu et al. 2023), similar to VLM2Scenes (Liao, Li, and Ye 2024), employs VLMs in distilling. Our approach achieved optimal results with 1% data fine-tuning on nuScenes and SemanticKITTI, reaching 51.75% mIoU and 48.14% mIoU, respectively, demonstrating a respective improvement of +21.45% mIoU and +8.64% mIoU versus random initialization. Compared to the previously best results, AFOV exhibited enhancements of +4.16% mIoU and +0.77% mIoU, respectively. Remarkably, the results of the fully supervised linear probing task on nuScenes reached 56.35% mIoU, displaying an improvement of +4.81% mIoU.
Ablation Study
@@
We conducted a series of ablation experiments on nuScenes. The ablation targets included different teacher models, TMP, AFI, etc. The results obtained validate the effectiveness of our designs, especially TMP and AFI. Please refer to Appendix B for details of the ablation study.
Conclusion